Let {an} ∞n=1 be a non-decreasing (resp. non-increasing) sequence which converges to a. Then prove that an ≤ a (resp. a ≤ an) for every n ∈ N.



1
Expert's answer
2021-02-12T18:09:41-0500

Let "\{a_n\}_{n=1}^{\infty}" be a non-decreasing sequence which converges to "a". Then prove that "a_n \u2264 a" for every "n \in\mathbb N."

Let us prove by contraposition. Suppose that "a_k>a" for some "k\in\mathbb N." Therefore, "a_k-a>0". Let "\epsilon =\frac{a_k-a}{2}". Since "\{a_n\}_{n=1}^{\infty}" be a convergent sequence with limit "a", there exist "M\in\mathbb N" such that "|a_n-a|<\epsilon" for every "n\ge M". Then "-\epsilon< a_n-a<\epsilon" for every "n\ge M". It follows that "a_n<\epsilon + a=\frac{a_k-a}{2}+a=\frac{a_k+a}{2}<\frac{a_k+a_k}{2}=a_k" for every "n\ge M". Let "N\ge \max\{k, M\}". Consequently, "a_n<a_k" for every "n\ge N", and we have a contradiction with the fact that "\{a_n\}_{n=1}^{\infty}" is a non-decreasing sequence.



Let "\{a_n\}_{n=1}^{\infty}" be a non-increasing sequence which converges to "a". Then prove that "a \u2264 a_n" for every "n \in\mathbb N."

Let us prove by contraposition. Suppose that "a>a_k" for some "k\in\mathbb N." Therefore, "a-a_k>0". Let "\epsilon =\frac{a-a_k}{2}". Since "\{a_n\}_{n=1}^{\infty}" be a convergent sequence with limit "a", there exist "M\in\mathbb N" such that "|a_n-a|<\epsilon" for every "n\ge M". Then "-\epsilon< a_n-a<\epsilon" for every "n\ge M". It follows that "a_n>a-\epsilon=a-\frac{a-a_k}{2}=\frac{a+a_k}{2}>\frac{a_k+a_k}{2}=a_k" for every "n\ge M". Let "N\ge \max\{k, M\}". Consequently, "a_n>a_k" for every "n\ge N", and we have a contradiction with the fact that "\{a_n\}_{n=1}^{\infty}" is a non-increasing sequence.